### v2

# Berezin transform of Toeplitz operators on Bergman space with Békollé and Bonami weights

Abstract.
In this paper, we obtain some interesting reproducing kernel estimates and some Carleson properties that play an important role. We completely characterized every case of the bounded and compact Toeplitz operators on the weighted Bergman spaces with Békollé-Bonami weights in terms of Berezin transforms.
Key words and phrases:
Békollé weights, Carleson measures, Toeplitz operators.
2010 Mathematics Subject Classification:
Primary: 47B35; Secondary: 32A36.
1. Introduction and Results
Let $\mathbb{C}$ be the complex plane, and for $r>0$, define the Euclidean open disk centered at the origin with radius $r$ as $D(0,r):=\{z\in\mathbb{C}:|z|<r\}$. The unit disc is defined as $\mathbb{D}:=D(0,1)$. Given a positive measure $\mu$ on $\mathbb{D}$ and $p>0$, the Lebesgue space $L^{p}(\mu)$ consists of all functions $f$ defined on $\mathbb{D}$ satisfying
$\|f\|_{L^{p}(\mu)}:=\left[\int_{\mathbb{D}}|f(z)|^{p}d\mu(z)\right]^{1/p}<\infty.$
For $p\geq 1$, $\|\cdot\|_{L^{p}(\mu)}$ defines a norm under which $L^{p}(\mu)$ forms a Banach space.
Let $dA$ denote the Lebesgue area measure on $\mathbb{D}$. If $u:\mathbb{D}\to\mathbb{R}^{+}$ is a locally integrable weight function (i.e., $u\in L^{1}_{loc}\left(dA\right)$), the weighted Lebesgue space $L^{p}(u)$ consists of all functions $f$ defined on $\mathbb{D}$ satisfying
$\|f\|_{L^{p}(u)}:=\left(\int_{\mathbb{D}}|f(z)|^{p}u(z)dA(z)\right)^{1/p}<\infty.$
The Bergman space $A^{p}(u)=A^{p}_{u}$ is defined as the closed subspace of analytic functions in $L^{p}(u)$ with $L^{p}(u)$-norm. When $u\equiv 1$, we write $A^{p}=A^{p}(1)$ for short. For the classical case $A^{2},$ the reproducing kernel is given by
$K_{w}(z)=\frac{1}{(1-\bar{w}z)^{2}}\text{ for }w,z\in\mathbb{D}.$
The following notations will be used throughout this paper. For a weight $u$ and a measurable set $E\subset\mathbb{D}$, we define $u(E)=\int_{E}udA$ and $A(E)=\int_{E}dA$. And
$\langle f\rangle^{d\mu}_{E}:=\frac{\int_{E}f(z)\textrm{d}\mu(z)}{\mu(E)},$
denotes the average of an integrable function $f$ over $E$ with respect to measure $\mu$.
The Bergman projection $P$ is defined by
$Pf(z)=\int_{\mathbb{D}}\frac{f(w)}{(1-\bar{w}z)^{2}}dA(w),\quad z\in\mathbb{D}.$
A classical problem in the theory of weighted Bergman spaces is to characterize the weights $u$, for which the Bergman projection $P$ is a bounded orthogonal projection from $L^{p}(u)$ to $A^{p}(u)$. This problem was solved by Békollé and Bonami ([1, 2]), they found that these weights are precisely $u\in B_{p},$ which is defined below.
The $B_{p}$ condition. Let $S(a)$ be the set
$S(a)=\left\{\frac{a-z}{1-\bar{a}z}:\mathrm{Re}(\bar{a}z)\leq 0\right\}.$
A weight $u$ satisfies the $B_{p}$ condition, or $u\in B_{p}$, if
$u]_{B_{p}}:={\sup_{a\in\mathbb{D}}}\langle u\rangle^{dA}_{S(a)}\left(\langle u^{-p^{\prime}/p}\rangle^{dA}_{S(a)}\right)^{p-1}\lesssim 1.$
Recently, the sharp estimates for the $L^{p}$-continuity of the Bergman projection are investigated in [11] and [12] respectively.
Toeplitz operators. The Bergman space $A^{2}(u)$ is endowed with the inner product
$\langle f,g\rangle_{A^{2}(u)}=\int_{\mathbb{D}}f(w)\overline{g(w)}u(w)dA(w),\quad f,g\in A^{2}(u),$
and its reproducing kernel is denoted by $K(z,w)$. Given a finite positive Borel measure $\mu$ on $\mathbb{D}$, the associated Toeplitz operator $T_{\mu}$ on $A^{2}(u)$ is defined as
$T_{\mu}f(z):=\int_{\mathbb{D}}f(w)K(z,w)d\mu(w),\quad z\in\mathbb{D}.$
This operator is well-defined when $\mu$ satisfies the condition
(1.1) $\int_{\mathbb{D}}|K(\xi,z)|^{2}d\mu(\xi)<\infty,\quad z\in\mathbb{D}.$
The pseudo-hyperbolic metric $d:\mathbb{D}\times\mathbb{D}\to[0,1)$ is defined as
$d(z,w)=\left|\frac{z-w}{1-\bar{w}z}\right|.$
For $z\in\mathbb{D}$ and $r>0$, the pseudo-hyperbolic disk centered at $z$ with radius $r$ is given by
(2.1) ${\Delta(z,r)}:=\{w\in\mathbb{D}:d(z,w)<r\}.$
and its volume is $\operatorname{vol}({\Delta(z,r)})=4\pi\frac{r^{2}}{1-r^{2}}$.
Given a finite positive Borel measure $\mu$ on $\mathbb{D}$ and $r>0$, the weighted average function $\widehat{\mu_{r}}$ is defined as
$\widehat{\mu_{r}}(z)=\frac{\mu(\Delta(z,r))}{u(\Delta(z,r))},\quad z\in\mathbb{D},$
where $u(\Delta(z,r))$ denotes the integral of the weight $u$ over the pseudo-hyperbolic disk $\Delta(z,r)$.
For $0<p<\infty$, the normalized reproducing kernel of $A^{p}_{u}$ is defined by $k_{p,z}(w):=K(w,z)/\|K(w,z)\|_{A^{p}_{u}}$. In particular, when $p=2$, we write $k_{z}(w):=k_{2,z}(w)$. The Berezin tranform is an important tool to study some proprieties of Toeplitz operators. The Berezin transform of the Toeplitz operator $T_{\mu}$ is given by
$\widetilde{\mu}(z):=\langle T_{\mu}k_{z},k_{z}\rangle_{A^{2}(u)},\quad z\in\mathbb{D},$
where $k_{z}(w):=K(w,z)/\|K(\cdot,z)\|_{A^{2}(u)}$ is the normalized Bergman kernel.
By direct computation one has
(1.2) $\langle T_{\mu}f,g\rangle_{A^{2}(u)}=\langle f,g\rangle_{L^{2}(d\mu)},$
which yields the integral representation of the Berezin transform
$\widetilde{\mu}(z)=\int_{\mathbb{D}}|k_{z}(w)|^{2}d\mu(w),\quad z\in\mathbb{D}.$
More generally, for $t>0$, the $t$-Berezin transform of $\mu$ is defined by
$\widetilde{\mu_{t}}(z)=\int_{\mathbb{D}}|k_{t,z}(w)|^{t}d\mu(w),\quad z\in\mathbb{D}.$
Our main propose is to obtain characterizations of some properties of Toeplitz operator $T_{\mu}:A_{u}^{q}\to A_{u}^{p}$ such as the boundedness, compactness and Schatten classmembership in terms of its generalized t-Berezin transform and average function, and another condition related to the geometric characterization of the symbol $\mu$ to be Carleson measure for $A^{s}_{u},$ for every possible cases of $s,p,q\in(0,\infty).$ The first two results give some equivalent necessary and sufficient conditions for the boundedness and compactness, respectively, of the operator $T_{\mu}$ acting between different Békollé and Bonami Bergman spaces $A_{u}^{q}(\mathbb{D})$ and $A_{u}^{q}(\mathbb{D}),$ for $0<p\leq q<\infty.$
Theorem 1.1.
Let $p_{0}>1$ and $u\in B_{p_{0}}$. When $0<p\leq q<\infty$, following assertions are equivalent:
(i):
The Toeplitz operator $T_{\mu}:A_{u}^{p}(\mathbb{D})\to A_{u}^{q}(\mathbb{D})$ is bounded;
(ii):
$\frac{\widetilde{\mu}_{t}(z)}{u(\Delta(z,r))^{\frac{1}{p}-\frac{1}{q}}}\in L^{\infty}$;
(iii):
$\frac{\widehat{\mu_{r}}(z)}{u(\Delta(z,r))^{\frac{1}{p}-\frac{1}{q}}}\in L^{\infty}$;
(iv):
$\mu$ is a $\frac{s(p+q^{\prime})}{pq^{\prime}}$-Carleson measure for $A_{u}^{s}$, where $\frac{1}{q}+\frac{1}{q^{\prime}}=1$ and $s\in(0,\infty)$.
Furthermore,
$\|T_{\mu}\|\simeq\left\|\frac{\widetilde{\mu}_{t}(z)}{u(\Delta(z,r))^{\frac{1}{p}-\frac{1}{q}}}\right\|_{L^{\infty}}\simeq\left\|\frac{\widehat{\mu_{r}}(z)}{u(\Delta(z,r))^{\frac{1}{p}-\frac{1}{q}}}\right\|_{L^{\infty}}.$
Theorem 1.2.
Let $p_{0}>1$ and $u\in B_{p_{0}}$. When $0<p\leq q<\infty$, following assertions are equivalent:
(i):
The Toeplitz operator $T_{\mu}:A_{u}^{p}(\mathbb{D})\to A_{u}^{q}(\mathbb{D})$ is compact;
(ii):
$\lim_{|z|\to 1}\frac{\widetilde{\mu}_{t}(z)}{u(\Delta(z,r))^{\frac{1}{p}-\frac{1}{q}}}=0$;
(iii):
$\lim_{|z|\to 1}\frac{\widehat{\mu_{r}}(z)}{u(\Delta(z,r))^{\frac{1}{p}-\frac{1}{q}}}=0$;
(iv):
$\mu$ is a $\frac{s(p+q^{\prime})}{pq^{\prime}}$-vanishing Carleson measure for $A_{u}^{s}$, where $\frac{1}{q}+\frac{1}{q^{\prime}}=1$ and $s\in(0,\infty)$.
The following theorem gives some equivalent necessary and sufficient conditions for the boundedness and compactness of Toeplitz operator $T_{\mu},$ for $0<q<p<\infty.$
Theorem 1.3.
Let $p_{0}>1$ and $u\in B_{p_{0}}$. When $0<q<p<\infty$, following assertion are equivalent:
(i):
The Toeplitz operator $T_{\mu}:A_{u}^{p}(\mathbb{D})\to A_{u}^{q}(\mathbb{D})$ is compact;
(ii):
The Toeplitz operator $T_{\mu}:A_{u}^{p}(\mathbb{D})\to A_{u}^{q}(\mathbb{D})$ is bounded;
(iii):
$\widetilde{\mu_{t}}\in L^{\frac{pq}{p-q}}$;
(iv):
$\widehat{\mu_{r}}\in L^{\frac{pq}{p-q}}$;
(v):
$\mu$ is a $(p+1-\frac{p}{q})$-Carleson measure for $A_{u}^{p}$;
(vi):
$\mu$ is a vanishing $(p+1-\frac{p}{q})$-Carleson measure for $A_{u}^{p}$.
Moreover, in the last results of this paper, we characterize the essential norm of $T_{\mu}$ and its Schatten class membership.
Theorem 1.4.
Let $\mu$ a finite Borel measure on $\mathbb{D}$. Suppose that $T_{\mu}:A^{p}_{u}\to A^{q}_{u}$ is a bounded operator. For any $t>0$ and $r>0$, we have
$\|T_{\mu}\|_{e}\simeq\limsup_{|z|\to 1}\widetilde{\mu_{t}}(z)u(\Delta(z,r))^{\frac{q-p}{pq}}\simeq\limsup_{|z|\to 1}\widehat{\mu_{r}}(z)u(\Delta(z,r))^{\frac{q-p}{pq}}.$
Theorem 1.5.
Let $h:\mathbb{R}^{+}\to\mathbb{R}^{+}$ a continuous increasing convex function. Let $\mu$ a positive Borel measure on $\mathbb{C}^{n}$ such that the Toeplitz operator $T_{\mu}:A_{u}^{2}\to A_{u}^{2}$ is compact. Then $T_{\mu}\in S_{h}(A_{u}^{2})$ if and only if there is a constant $C$ such that
$\int_{\mathbb{C}^{n}}h(C\widetilde{\mu}_{2}(z))\frac{u(\Delta(z,r))}{(1-|z|)^{4}}u(z)dA(z)<\infty.$
Throughout the paper, we will use following notations:
  • $Q_{1}\lesssim Q_{2}$ means that there is a constant $C>0$ (independent of the key variable(s)) such that $Q_{1}\leq CQ_{2}$;
  • $Q_{1}\simeq Q_{2}$ if both $Q_{1}\lesssim Q_{2}$ and $Q_{2}\lesssim Q_{1}$;
  • For $p>1$, $p^{\prime}$ is the conjugate exponent of $p$, i.e., $\frac{1}{p}+\frac{1}{p^{\prime}}=1$.
2. preliminaries and basic properties
In this section we review a number of preliminary results related to the weight class, characterizations of Carleson measures and the Bergman kernels in the corresponding weighted spaces, which lead us to prove our main results.
The pseudo-hyperbolic metric satisfies the following strong triangle inequality
(2.1) $d(z,w)\leq\frac{d(z,\zeta)+d(\zeta,w)}{1+d(z,\zeta)d(\zeta,w)},\quad\forall z,w,\zeta\in\mathbb{D}.$
For $0<r<1$, if $z,w\in\mathbb{D}$ satisfy $d(z,w)<r$, then the following approximations hold,
(2.2) $1-|z|\simeq 1-|w|\simeq|1-\bar{w}z|.$
Moreover, for any $\zeta\in\mathbb{D}$, we have
(2.3) $\left|\frac{1-\bar{\zeta}z}{1-\bar{\zeta}w}\right|\simeq 1,$
where the constants involved depend only on $r$. Furthermore, we define a modified pseudohyperbolic disk which is compareable with $\Delta(z,r)$(2.1),
$\Delta^{\prime}(z,r):=\left\{w\in\mathbb{D}:\frac{|z-w|}{1-|w|}<r\right\}.$
We will also use the following class of weights which is denoted by condition $C_{p}$. For $p>1$, a positive locally integrable weight $u$ belongs to the class $C_{p}$ (or satisfies condition $C_{p}$) if
$[u]_{C_{p}}:={\sup_{z\in\mathbb{D}}}\langle u\rangle^{\textrm{d}A}_{\Delta(z,r)}\left(\langle u^{-p^{\prime}/p}\rangle^{\textrm{d}A}_{\Delta(z,r)}\right)^{p-1}\lesssim 1.$
It is known that this class of weights is obtained for any $r\in(0,1)$ and $B_{p}\subset C_{p}$. To see this, we note that for a given $r$, there is a $a^{\prime}\in\mathbb{D}$ such that $\Delta(a,r)\subset S(a^{\prime})$ with comparable volumes, see more details in [7].
For $p>0$, the point evaluations on $A^{p}(u)$ are bounded linear functionals, as one can see from the following lemma.
Lemma 2.1 (Theorem 3.1 in [7]).
If $p_{0}>1$, $p>0$, $0<r<1$ and a weight $u\in C_{p_{0}}$, we have
$|f(z)|^{p}\leq Cu(\Delta(z,r))^{-1}\int_{\Delta(z,r)}|f(w)|^{p}u(w)\textrm{d}A(w),$
where the constant $C>0$ depends on $r,p$.
Within the Békollé-Bonami class of weights, Bergman metric balls with centers sufficiently close in the pseudo-hyperbolic metric have comparable weighted areas.
Lemma 2.2 (Lemma 2.2 in [5]).
Suppose $u\in C_{p}$ for some $p>1$. Let $t,s\in(0,1)$, and $z,w\in\mathbb{D}$ with $d(z,w)<r$ for some $r>0$. Then we have
$u(\Delta({z},{t}))\simeq u(\Delta({w},s)),$
where the constant is independent of $z$ and $w$.
Similarly, if $u\in B_{p_{0}}$, it should be noted that
(2.4) $u(\Delta(a,r))\simeq u(S(a^{\prime})),$
whenever $\Delta(a,r)\subset S(a^{\prime})$ with comparable areas. For further details, see [7] and Lemma 5.23 in [16].
Our method relies on a popular decomposition of $\mathbb{D}$ which is used repeatedly in many papers, e.g., Theorem 1 in [15]. We will use the following covering lemma.
Lemma 2.3 (Theorem 2.23 in [16]).
There exists a positive $N$ such that for any $0<r\leq 1$ we can find a sequence $\{a_{k}\}$ in $\mathbb{D}$ with the following properties.
(1):
$\mathbb{D}=\cup_{k}\Delta({a_{k}},{r})$;
(2):
The set $\Delta({a_{k}},{r/4})$ are mutually disjoint;
(3):
Each point $z\in\mathbb{D}$ belongs to at most $N$ of the sets $\Delta({a_{k}},{2r})$.
2.1. Carleson measures
A positive Borel measure $\mu$ on $\mathbb{D}$ is a qq-(vanishing) Carleson measure for $A^{p}(u)$ if the embedding $I:A^{p}(u)\to L^{q}(\textrm{d}\mu)$ is (compact) bounded. The following two lemma give the characterizations of $q$- (vanishing) Carleson measure for $A^{p}(u),$ when $0<p\leq q$ and $u$ belongs to our class of weights.
To prove our main results we need some characterizations of the bounded and compact embedding operator $I,$ whenever $0<p,q<\infty,$ which are presented in the following lemmas.
Lemma 2.4 (Theorem 3.1 in [13]).
Suppose $q\geq p>0$, $p_{0}>1$ and $0<r<1$. Let $u\in B_{p_{0}}$ be a weight and $\mu$ is a positive Borel measure on $\mathbb{D}$. Then the following conditions are equivalent.
(a):
The embedding $I:A^{p}(u)\to L^{q}(\textrm{d}\mu)$ is bounded, that is
$\left(\int_{\mathbb{D}}|f(z)|^{q}\textrm{d}\mu(z)\right)^{1/q}\lesssim\left(\int_{\mathbb{D}}|f(z)|^{p}u(z)\textrm{d}A(z)\right)^{1/p}$
for every analytic function $f$ on $\mathbb{D}$;
(b):
$\mu(S(a))\lesssim u(S(a))^{q/p}$ for every $a\in\mathbb{D}$;
(c):
There is an $r>0$ such that $\mu({\Delta(a,r)})\lesssim u({\Delta(a,r)})^{q/p}$ for every $a\in\mathbb{D}$;
(d):
There is an $r>0$ such that $\mu({\Delta(a_{k},r)})\lesssim u({\Delta(a_{k},r)})^{q/p}$ for the sequence $\{a_{k}\}_{k}$ described in Lemma 2.3;
(e):
Denoted by
$g_{w}^{s}(z)=\frac{1}{u(\Delta(w,r))^{q/p}}\left(\frac{1-|w|^{2}}{1-z\bar{w}}\right)^{s},$
For any $s\geq 2p_{0}/p$
(2.5) $\|g_{w}^{s}\|_{L^{q}(\textrm{d}\mu)}^{q}=\int_{\mathbb{D}}\left|\frac{1-|w|^{2}}{1-z\bar{w}}\right|^{qs}u({\Delta(w,r)})^{-q/p}\textrm{d}\mu(z)\lesssim 1.$
Furthermore, the “geometric norm” of the measure $\mu$, the $L^{q}(\textrm{d}\mu)$ norm of $g^{s}_{w}$ and the operator norm of the embedding are comparable:
${\sup_{z\in\mathbb{D}}}\widehat{\mu_{r}}(z):={\sup_{z\in\mathbb{D}}}\frac{\mu({\Delta(z,r)})}{u({\Delta(z,r)})^{q/p}}\simeq{\sup_{w\in\mathbb{D}}}\|g^{s}_{w}\|^{q}_{L^{q}(\textrm{d}\mu)}\simeq\|I\|^{q}_{A^{p}(u)\to L^{q}(\textrm{d}\mu)}.$
The equivalence (a) and (c) was proved by Constantin in [5].
Lemma 2.5 (Theorem 3.3 in [13]).
Suppose $q\geq p>0$, $p_{0}>1$ and $0<r<1$. Let $u\in B_{p_{0}}$ be a weight and $\mu$ is a positive Borel measure on $\mathbb{D}$. Then $I_{d}:A_{u}^{p}\to L^{q}_{\mu}$ is compact if and only if:
$\lim_{|z|\to 1^{-}}\frac{\mu(S(a))}{(u(S(a)))^{q/p}}=0.$
2.2. Kernel Estimates
Recall that $K(w,z)=K_{w}(z)$ denotes the reproducing kernels of $A_{u}^{2}$, and $k_{w}(z)=\frac{|K_{w}(z)|}{\|K_{w}\|_{A_{u}^{2}}}$ denotes the normalized reproducing kernel of $A_{u}^{2}$. In what follows we will introduce a $L^{p}$-estimate of $K_{w}(z)$ and also a pointwise estimate of $|k_{w}(z)|^{2}$.
And by similar arguments in Lemma 3.1 of [4], when taking $\beta=\gamma=0$ in it we get following Lemma.
Lemma 2.6.
Let $0<p<\infty$, $p_{0}>1$ and $u\in B_{p_{0}}$. Then we have the following estimate:
(2.6) $\|K_{w}\|_{A_{u}^{p}}\simeq\frac{u(\Delta(w,r))^{\frac{1}{p}}}{(1-|w|)^{2}}.$
And recall thet the normalized reproducing kernel $k_{p,w}=\frac{K_{w}}{\|K_{w}\|_{A_{u}^{p}}}$, by Lemma 2.1 and Lemma 2.6 above, we have
(2.7) $|k_{p,w}(z)|\lesssim u(\Delta(w,r))^{-\frac{1}{p}},\quad\textrm{for any}\,\,\,z,w\in\mathbb{D}.$
Next, the following lemma shows the estimates of the reproducing kernel functions on the diagonal.
Lemma 2.7 (Lemma 4.1 in [5]).
Suppose $p_{0}>1$ and $u\in B_{p_{0}}$. Let $K(z,w)$ be the Bergman kernel in $A^{2}(u)$ and $r\in(0,1)$. Then we have the following estimate
(2.8) $K(z,z)\simeq u(\Delta(z,r))^{-1},\quad z\in\mathbb{D},$
where the constant involved is independent of $z\in\mathbb{D}$.
Now, the following lemma gives the pointwise estimates of $k_{w}(z),$ when $z$ and $w$ are closed.
Lemma 2.8 (Lemma 2.7 in [14]).
Suppose $p_{0}>1$ and $u\in B_{p_{0}}$. There is a sufficient small $\delta\in(0,1)$, such that
$|k_{w}(z)|^{2}\simeq K(z,z),$
whenever $z\in\Delta(w,\delta)$.
The remain Lemmas of this subsection show the relationship between Berezin transform $\tilde{\mu}_{t}(z)$ and the average function $\widehat{\mu_{r}}(z),$ for any $r,t>0.$
Lemma 2.9.
For a positive Borel measure $\mu$ on $\mathbb{D}$ and $r>0$. If $\mu$ satisfies the condition
(2.9) $\int_{\mathbb{D}}|K(\xi,z)|^{2}\textrm{d}\mu(\xi)<\infty.$
Then,
$\tilde{\mu}(z)\simeq\widehat{\mu_{r}}(z).$
Proof.
Since $0<r\leq\delta$, we use Lemma 2.8, Lemma 2.7 and Lemma 2.2 to see that
(2.10) $\tilde{\mu}(z)\geq\int_{\Delta(z,r)}|k_{z}(\zeta)|^{2}\textrm{d}\mu(\zeta)\simeq\int_{\Delta(z,r)}K(\zeta,\zeta)\textrm{d}\mu(\zeta)$
(2.11) $\simeq$ $\int_{\Delta(z,r)}\frac{1}{u(\Delta(\zeta,r))}\,\textrm{d}\mu(\zeta)\simeq\frac{1}{u(\Delta(z,r))}\int_{\Delta(z,r)}\,\textrm{d}\mu(\zeta)=\widehat{\mu_{r}}(z).$
Conversely, by Lemma 2.1, one has that
$|k_{z}(w)|^{2}\lesssim\frac{1}{u(\Delta(w,r))}\int_{\Delta(w,r)}|k_{z}(\zeta)|^{2}u(\zeta)\,\textrm{d}A(\zeta),$
where $z,w\in\mathbb{D}$ and $r>0$. Let $\{a_{j}\}$ and $r>0$ be chosen as in Lemma 2.3, using Lemma 2.2, Fubini’s Theorem and Lemma 2.3, we can conduct the following computation,
$\tilde{\mu}(z)\leq\sum\limits_{j=1}^{\infty}\int_{\Delta(a_{j},r)}|k_{z}(w)|^{2}\,\textrm{d}\mu(w)$
$\lesssim$ $\sum\limits_{j=1}^{\infty}\int_{\Delta(a_{j},r)}\frac{1}{u(\Delta(w,r))}\int_{\Delta(w,r)}|k_{z}(\zeta)|^{2}u(\zeta)\,\textrm{d}A(\zeta)\,\textrm{d}\mu(w)$
$\lesssim$ ${\sup_{j}}\frac{\mu(\Delta(a_{j},r))}{u(\Delta(a_{j},r))}\sum\limits_{j=1}^{\infty}\int_{\Delta(a_{j},2r)}|k_{z}(\zeta)|^{2}u(\zeta)\,\textrm{d}A(\zeta)$
$\lesssim$ ${\sup_{z\in\mathbb{D}}}\widehat{\mu_{r}}(z),$
where the last inequality follows from Lemma 2.3. The proof is complete. ∎
Lemma 2.10.
For a positive Borel measure $\mu$ on $\mathbb{D}$ and $r>0$. If $\mu$ satisfies the condition
(2.12) $\int_{\mathbb{D}}|K(\xi,z)|^{p}\textrm{d}\mu(\xi)<\infty.$
for $0<p<\infty$. Then, The following assertion are equivalent:
(i):
$\widetilde{\mu_{t}}\in L^{p}$ for any $t>0$;
(ii):
$\widehat{\mu_{r}}\in L^{p}$.
Furthermore,
$\tilde{\mu}_{t}(z)\simeq\widehat{\mu_{r}}(z).$
Proof.
[…]
$\tilde{\mu}_{t}(z)$ $\geq$ $\int_{\Delta(z,r)}|k_{t,z}(w)|^{t}d\mu(w)$
$=$ $\int_{\Delta(z,r)}\left|\frac{k_{z}(w)\|K_{z}\|_{A_{u}^{2}}}{\|K_{z}\|_{A_{u}^{t}}}\right|^{t}d\mu(w)$
$\simeq$ $u(\Delta(z,r))^{\frac{t}{2}-1}\int_{\Delta(z,r)}\left|{K(w,w)}\right|^{\frac{t}{2}}d\mu(w)$
$\simeq$ $\widehat{\mu_{r}}(z).$
[…]
$\tilde{\mu}_{t}(z)\leq\sum\limits_{j=1}^{\infty}\int_{\Delta(a_{j},r)}|k_{t,z}(w)|^{t}\,\textrm{d}\mu(w)$
$\lesssim$ $\sum\limits_{j=1}^{\infty}\int_{\Delta(a_{j},r)}\frac{1}{u(\Delta(w,r))}\int_{\Delta(w,r)}|k_{t,z}(\zeta)|^{t}u(\zeta)\,\textrm{d}A(\zeta)\textrm{d}\mu(w)$
$\lesssim$ ${\sup_{j}}\frac{\mu(\Delta(a_{j},r))}{u(\Delta(a_{j},r))}\sum\limits_{j=1}^{\infty}\int_{\Delta(a_{j},2r)}|k_{t,z}(\zeta)|^{t}u(\zeta)\,\textrm{d}A(\zeta)$
$\lesssim$ ${\sup_{z\in\mathbb{D}}}\widehat{\mu_{r}}(z)\,\|k_{t,z}\|^{t}_{A^{t}_{u}}={\sup_{z\in\mathbb{D}}}\widehat{\mu_{r}}(z).$
The following corollary follows directly from Lemma 2.9 and Lemma 2.10.
Corollary 2.11.
For a positive Borel measure $\mu$ on $\mathbb{D}$, $0<p<\infty$ and $r>0$. If $\mu$ satisfies the condition
(2.13) $\int_{\mathbb{D}}|K(\xi,z)|^{p}\textrm{d}\mu(\xi)<\infty.$
Then the following assertion are equivalent:
(i):
$\widetilde{\mu_{t}}\in L^{p}$ for any $t>0$;
(ii):
$\widetilde{\mu}\in L^{p}$.
Furthermore,
$\tilde{\mu}(z)\simeq\tilde{\mu}_{t}(z)\simeq\widehat{\mu_{r}}(z).$
2.3. Duality
In this part, we establish a duality relation for weights in the $B_{p_{0}}$ class. As normal, we define $p^{\prime}$ as the conjugate for a $p>1$, characterised by the relation $\frac{1}{p}+\frac{1}{p^{\prime}}=1$.
Lemma 2.12.
Suppose $1<p_{0}$ and $u\in B_{p_{0}}$. For $1<p<\infty$, The dual of $A_{u}^{p}$ can be identified with $A_{u}^{p^{\prime}}$ under the pairing
(2.14) $\langle f,g\rangle_{u}=\int_{\mathbb{D}}f(z)\overline{g(z)}u(z)\textrm{d}A(z).$
Proof.
For all $f\in A_{u}^{p^{\prime}}$ and $g\in A_{u}^{p}$, denote $\Lambda_{g}(f)=\langle f,g\rangle_{u}$. Then, using Hölder’s inequality, we have
$|\Lambda_{g}(f)|=|\langle f,g\rangle_{u}|\leq\int_{\mathbb{D}}|f(z)||g(z)|u(z)dA(z)\leq\|f\|_{A_{u}^{p^{\prime}}}\|g\|_{A_{u}^{p}},$
so $\Lambda_{g}\in(A_{u}^{p^{\prime}})^{*}$, and $\|\Lambda_{g}\|\leq\|g\|_{A_{u}^{p}}.$
By the Hahn-Banach theorem, every bounded linear operator $T\in(A_{u}^{p^{\prime}})^{*}$ can be extended to a bounded linear functional $\widetilde{T}$ on $L_{u}^{p^{\prime}}$. Since $(A_{u}^{p^{\prime}})^{*}$ is isometrically isomorphic to $L_{u}^{p}$, there exists an $h\in L_{u}^{p}$ such that $T(f)=\langle f,h\rangle_{u}$ for all $f\in A_{u}^{p^{\prime}}$, and $\|\widetilde{T}\|=\|h\|_{L_{u}^{p}}$. Since the Bergman projection $P_{u}:L^{p}_{u}\to A^{p}_{u}$ is bounded ([1, 2]) and then using Fubini’s theorem, for any $f\in A_{u}^{p^{\prime}}$ we have
$T(f)=T(P_{u}f)=\langle f,P_{u}h\rangle_{u}=\Lambda_{g}(f),$
where $g=P_{u}h$. It follows that $g\in A_{u}^{p}$ and satisfies $\|g\|_{A_{i}^{p}}\leq\|P_{u}\|\|h\|_{L_{u}^{p}}=\|P_{u}\|\|T\|$. This completes the proof. ∎
Then we can prove the following lemma, which is essential for characterizing compactness properties of operators acting on Békollé and Bonami Bergman spaces. Chacón [3] proved a version for $k_{2,w}$ case, we show it below for convenient.
Lemma 2.13 (Propositon 19 in [3]).
Let $p_{0}>1$. If $u\in B_{p_{0}}$, then the normalized kernel function $k_{p,w}$ is weakly converge to zero as $|w|\to 1^{-}$ in $A^{2}_{u}$.
Then we generalize Lemma 2.13 to arbitrary $p$-case.
Lemma 2.14.
Let $p_{0}>1$, $1<p<\infty$. If $u\in B_{p_{0}}$, then the normalized kernel function $k_{p,w}$ is weakly converge to zero as $|w|\to 1^{-}$.
Proof.
By Lemma 2.12, for $1<p<\infty$, it suffices to show that
$|\langle k_{w},g\rangle_{u}|\to 0,\text{ as }|w|\to 1^{-},$
for all $g\in A_{u}^{p^{\prime}}$ and $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. By the definition of $k_{w}$ and reproducing kernel formula, we get
$|\langle k_{p,w},g\rangle_{u}|$ $=$ $\frac{1}{\|K_{w}\|_{A_{u}^{p}}}\Big|\int_{\mathbb{D}}K(z,w)\,g(z)\,u(z)\,dA(z)\Big|$
$=$ $\frac{|g(w)|}{\|K_{w}\|_{A_{u}^{p}}}.$
By the kernel norm estimate from Lemma 2.6 and Lemma 2.1, we get
$|\langle k_{p,w},g\rangle_{u}|\simeq\frac{|g(w)|(1-|w|)^{2}}{u(\Delta(w,r))^{\frac{1}{p}}},$
$|\langle k_{p,w},g\rangle_{u}|$ $\simeq$ $\frac{|g(w)|(1-|w|)^{2}}{u(\Delta(w,r))^{\frac{1}{p}}}$
$\lesssim$ $\frac{(1-|w|)^{2}}{u(\Delta(w,r))^{\frac{1}{p}}}\left(u(\Delta(w,r))^{-1}\int_{\Delta(w,r)}|g(z)|^{p^{\prime}}u(z)dA(z)\right)^{\frac{1}{p^{\prime}}}$
$\lesssim$ $\frac{(1-|w|)^{2}}{u(\Delta(w,r))}\|g\|_{A^{p^{\prime}}_{u}}$
$\simeq$ $\left|\langle k_{2,w},\frac{1}{u(\Delta(w,r))^{\frac{1}{2}}}\rangle_{u}\right|\|g\|_{A^{p^{\prime}}_{u}}$
then Lemma 2.13 implies that $k_{p,w}$ converges to zero weakly as $|w|\to 1^{-}.$
The following Proposition is a classical result and its proof is similar to one given by K. Zhu in [17, Theorem 1.14].
Proposition 2.15.
Let $p_{0}>1$ and $u\in B_{p_{0}}.$ A linear operator $T$ on $A^{2}(u)$ is compact if and only if $\|Tf_{n}\|_{A^{2}(u)}\longrightarrow 0,$ whenever $f_{n}\to 0$ weakly in $A^{2}(u).$
3. Proof of Theorem 1.1
Proof.
[…]
$|\widetilde{\mu}(z)|$ $=$ $\left|\langle T_{\mu}k_{z},k_{z}\rangle_{A^{2}_{u}}\right|\leq\|T_{\mu}k_{z}\|_{A^{q}_{u}}\|k_{z}\|_{A_{u}^{q^{\prime}}}$
$\leq$ $\|T_{\mu}\|\|k_{z}\|_{A_{u}^{p}}\|k_{z}\|_{A_{u}^{q^{\prime}}}$
$\leq$ $\|T_{\mu}\|\frac{u(\Delta(z,r))^{1/p}}{u(\Delta(z,r))^{1/2}}\frac{u(\Delta(z,r))^{1/q^{\prime}}}{u(\Delta(z,r))^{1/2}}$
$\simeq$ $\|T_{\mu}\|u(\Delta(z,r))^{\frac{1}{p}-\frac{1}{q}},$
which implies that
$\frac{|\widetilde{\mu}(z)|}{u(\Delta(z,r))^{\frac{1}{p}-\frac{1}{q}}}\lesssim\|T_{\mu}\|<\infty.$
For the case $q\leq 1$, by Lemma 2.8, inequality (2.7) and then Lemma 2.1, we have
$\Big|\frac{\widetilde{\mu}(z)}{u(\Delta(z,r))^{\frac{1}{p}-\frac{1}{q}}}\Big|$ $=$ $\Big|u(\Delta(z,r))^{\frac{1}{q}-\frac{1}{p}}\int_{\mathbb{D}}|k_{z}(w)|^{2}d\mu(w)\Big|$
$\simeq$ $\Big|u(\Delta(z,r))^{\frac{1}{q}-\frac{1}{p}}\int_{\mathbb{D}}\frac{|K(z,w)|^{2}}{\|K_{z}\|_{A_{u}^{2}}^{2}}d\mu(w)\Big|$
$\simeq$ $\Big|u(\Delta(z,r))^{\frac{1}{q}-\frac{1}{p}}\frac{\|K_{z}\|_{A_{u}^{p}}}{\|K_{z}\|_{A_{u}^{2}}^{2}}\int_{\mathbb{D}}K(z,w)\frac{\overline{K(z,w)}}{\|K_{z}\|_{A_{u}^{p}}}d\mu(w)\Big|$
$\simeq$ $\Big|u(\Delta(z,r))^{\frac{1}{q}}\int_{\mathbb{D}}K(z,w)k_{p,z}(w)\textrm{d}\mu(w)\Big|$
$\simeq$ $u(\Delta(z,r))^{\frac{1}{q}}|T_{\mu}k_{p,z}|$
$\lesssim$ $\left(\int_{\Delta(z,r)}|T_{\mu}k_{p,z}|^{q}u(\zeta)dA(\zeta)\right)^{\frac{1}{q}}$
$\lesssim$ $\|T_{\mu}k_{p,z}\|_{A_{u}^{q}}\leq\|T_{\mu}\|.$
Thus,
$\frac{|\widetilde{\mu}(z)|}{u(\Delta(z,r))^{\frac{1}{p}-\frac{1}{q}}}\in L^{\infty}.$
In addition, by Corollary 2.11, we have $\frac{|\widetilde{\mu}(z)|}{u(\Delta(z,r))^{\frac{1}{p}-\frac{1}{q}}}\simeq\frac{|\widetilde{\mu}_{t}(z)|}{u(\Delta(z,r))^{\frac{1}{p}-\frac{1}{q}}}\in L^{\infty}$, for any $t>0.$
[…]
$\mu(S)\lesssim u(S)^{\frac{p+q^{\prime}}{pq^{\prime}}},$
[…]
$\left(\int_{\mathbb{D}}|f(z)|^{\frac{p+q^{\prime}}{pq^{\prime}}}d\mu(z)\right)^{\frac{pq^{\prime}}{p+q^{\prime}}}$ $\lesssim\int_{\mathbb{D}}|f(z)|u(z)dA(z)$
$\int_{\mathbb{D}}|f(z)|^{\frac{p+q^{\prime}}{pq^{\prime}}}d\mu(z)$ $\lesssim\left(\int_{\mathbb{D}}|f(z)|u(z)dA(z)\right)^{\frac{p+q^{\prime}}{pq^{\prime}}}$
$\int_{\mathbb{D}}|h(z)|\,d\mu(z)$ $\lesssim\left(\int_{\mathbb{D}}|h(z)|^{{}^{\frac{pq^{\prime}}{p+q^{\prime}}}}u(z)dA(z)\right)^{\frac{p+q^{\prime}}{pq^{\prime}}},$
for any analytic function $f$ which could have zero points on $\mathbb{D}$, we consider $f_{\varepsilon}=f+\varepsilon,\varepsilon>0$, which is also analytic without zeros on $\mathbb{D}.$ By Fatou’s Lemma, and (3) and dominated convergence theorem, we get
$\int_{\mathbb{D}}|f|d\mu=\lim_{\varepsilon\to 0}\int_{\mathbb{D}}|f_{\varepsilon}|d\mu\leq\liminf_{\varepsilon\to 0}C\left(\int_{\mathbb{D}}|f_{\varepsilon}|^{\frac{pq^{\prime}}{p+q^{\prime}}}u\,dA\right)^{\frac{p+q^{\prime}}{pq^{\prime}}}\lesssim\left(\int_{\mathbb{D}}|f|^{\frac{pq^{\prime}}{p+q^{\prime}}}u\,dA\right)^{\frac{p+q^{\prime}}{pq^{\prime}}},$
which implies that $\mu$ is a 1-Carleson measure for $A_{u}^{\frac{pq^{\prime}}{p+q^{\prime}}}.$ This together with applying Hölder’s inequality, for any analytic function $f\in A^{p}_{u},$ we obtain
$\left|\langle T_{\mu}f,g\rangle_{A_{u}^{2}}\right|$ $\leq$ $\int_{\mathbb{D}}|f(z)g(z)|d\mu(z)$
$\lesssim$ $\left(\int_{\mathbb{D}}|f(z)g(z)|^{{}^{\frac{pq^{\prime}}{p+q^{\prime}}}}u(z)dA(z)\right)^{\frac{p+q^{\prime}}{pq^{\prime}}}$
$\lesssim$ $\|f\|_{A_{u}^{p}}\|g\|_{A^{q^{\prime}}_{u}},$
where $g\in A^{q^{\prime}}_{u}$ since duality (Lemma 2.12). It follows that $T_{\mu}:A^{p}_{u}\to A^{q}_{u}$ is bounded.
We now consider the case $q\leq 1$. Using some arguments like in the case $q>1$, we can show that $\mu$ is also a 1-Carleson measure for $A_{u}^{\frac{pq^{\prime}}{p+q^{\prime}}}$. Since $q\leq 1$, the triangle inequality $(a+b)^{q}\leq a^{q}+b^{q}$ holds for non-negative $a$ and $b$. For $f\in A_{u}^{p}$ and $\{z_{k}\}_{k}$ is a $r$-lattice as defined in Lemma 2.3. By Lemma 2.1, we obtain
$|T_{\mu}f(z)|^{q}$ $=$ $\left|\int_{\mathbb{D}}f(w)K(z,w)d\mu(w)\right|^{q}$
$\lesssim$ $\left|\left(\int_{\mathbb{D}}\left|f(w)K(z,w)\right|^{\frac{pq^{\prime}}{p+q^{\prime}}}u(w)dA(w)\right)^{\frac{p+q^{\prime}}{pq^{\prime}}}\right|^{q}$
$\lesssim$ $\sum_{k=1}^{\infty}\left|\left(\int_{\Delta(z_{k},r)}\left|f(w)K(z,w)\right|^{\frac{pq^{\prime}}{p+q^{\prime}}}u(w)dA(w)\right)^{\frac{p+q^{\prime}}{pq^{\prime}}}\right|^{q}$
$\lesssim$ $\sum_{k=1}^{\infty}\sup_{w\in\Delta(z_{k},r)}\left|f(w)K(z,w)\right|^{q}\left|\int_{\Delta(z_{k},r)}u(w)dA(w)\right|^{\frac{q(p+q^{\prime})}{pq^{\prime}}}$
$\simeq$ $\sum_{k=1}^{\infty}u(\Delta(z_{k},r))^{\frac{q(p+q^{\prime})}{pq^{\prime}}}\sup_{w\in\Delta(z_{k},r)}\left|f(w)K(z,w)\right|^{q}$
$\lesssim$ $\sum_{k=1}^{\infty}u(\Delta(z_{k},r))^{\frac{q(p+q^{\prime})}{pq^{\prime}}-1}\int_{\Delta(z_{k},r)}\left|f(\xi)\right|^{q}\left|K(z,\xi)\right|^{q}u(\xi)dA(\xi)$
$\simeq$ $\sum_{k=1}^{\infty}\int_{\Delta(z_{k},r)}u(\Delta(\xi,r))^{\frac{q(p+q^{\prime})}{pq^{\prime}}-1}\left|f(\xi)\right|^{q}\left|K(z,\xi)\right|^{q}u(\xi)dA(\xi)$
By Fubini’s theorem, Lemma 2.6 and Lemma 2.1, we have
$\|T_{\mu}f\|_{A_{u}^{q}}^{q}$ $\lesssim$ $\int_{\mathbb{D}}\sum_{k=1}^{\infty}\int_{\Delta(z_{k},r)}u(\Delta(\xi,r))^{\frac{q(p+q^{\prime})}{pq^{\prime}}-1}\left|f(\xi)\right|^{q}\left|K(z,\xi)\right|^{q}u(\xi)dA(\xi)u(z)dA(z)$
$\lesssim$ $\int_{\mathbb{D}}\int_{\mathbb{D}}u(\Delta(\xi,r))^{\frac{q(p+q^{\prime})}{pq^{\prime}}-1}\left|f(\xi)\right|^{q}\left|K(z,\xi)\right|^{q}u(\xi)dA(\xi)u(z)dA(z)$
$=$ $\int_{\mathbb{D}}u(\Delta(\xi,r))^{\frac{q(p+q^{\prime})}{pq^{\prime}}-1}\|K_{\xi}\|_{A_{u}^{q}}^{q}\left|f(\xi)\right|^{q}u(\xi)dA(\xi)$
$=$ $\int_{\mathbb{D}}u(\Delta(\xi,r))^{\frac{q(p+q^{\prime})}{pq^{\prime}}-1}\|K_{\xi}\|_{A_{u}^{q}}^{q}|f(\xi)|^{p}|f(\xi)|^{p\cdot\frac{q-p}{p}}u(\xi)dA(\xi)$
$\lesssim$ $\int_{\mathbb{D}}u(\Delta(\xi,r))^{\frac{q(p+q^{\prime})}{pq^{\prime}}-1}\|K_{\xi}\|_{A_{u}^{q}}^{q}|f(\xi)|^{p}$
$\cdot\left(u(\Delta(\xi,r))^{-1}\int_{u(\Delta(\xi,r))}|f(w)|^{p}u(w)dA(w)\right)^{\frac{q-p}{p}}u(\xi)dA(\xi)$
$\lesssim$ $\|f\|_{A_{u}^{p}}^{q-p}\int_{\mathbb{D}}u(\Delta(\xi,r))^{q-1}\|K_{\xi}\|_{A_{u}^{q}}^{q}|f(\xi)|^{p}u(\xi)dA(\xi)$
$\simeq$ $\|f\|_{A_{u}^{p}}^{q-p}\int_{\mathbb{D}}\frac{u(\Delta(\xi,r))^{q}}{(1-|\xi|)^{2q}}|f(\xi)|^{p}u(\xi)dA(\xi)$
$\simeq$ $\|f\|_{A_{u}^{p}}^{q-p}\int_{\mathbb{D}}\|K_{\xi}\|_{A_{u}^{1}}^{q}|f(\xi)|^{p}u(\xi)dA(\xi)$
$\leq$ $\sup_{\xi\in\mathbb{D}}\|K_{\xi}\|_{A_{u}^{1}}^{q}\|f\|_{A_{u}^{p}}^{q}$
which implies that $T_{\mu}$ is bounded. ∎
4. Proof of Theorem 1.2
Proof.
[…]
Which is a direct result of Lemma 2.10.
[…]
Which is a direct result of Lemma 2.5.
[…]
Suppose that $T_{\mu}$ is compact and consider the case $q>1$, then by Hölder’s inequality we have
$|\widetilde{\mu}(z)|=|\langle T_{\mu}k_{z},k_{z}\rangle|_{A_{u}^{2}}\leq\|T_{\mu}k_{z}\|_{A_{u}^{q}}\|k_{z}\|_{A_{u}^{q^{\prime}}}$
By Lemma 2.14, we have that $\|k_{z}\|_{A_{u}^{p}}\to 0$ as $|z|\to 1^{-}$. Since $T_{\mu}$ is compact, it follows that $\|T_{\mu}k_{z}\|_{A_{u}^{q}}\to 0$ as $|z|\to 1^{-}$, which implies that $|\widetilde{\mu}(z)|\to 0$ as $|z|\to 1^{-}$ .
For the case $q\leq 1$, we use a similar argument in the proof of Theorem 1.1 to obtain
$\frac{|\widetilde{\mu}(z)|}{u(\Delta(z,r))^{\frac{1}{p}-\frac{1}{q}}}\lesssim\|T_{\mu}k_{p,z}\|_{A_{u}^{q}}.$
Again, by Lemma 2.14 and using arguments analogous to those in the case $q>1$, we have $|\widetilde{\mu}(z)|\to 0$ as $|z|\to 1^{-}$ .
[…]
Suppose that $\mu$ is a $\frac{s(p+q^{\prime})}{pq^{\prime}}$-vanishing Carleson measure for $A_{u}^{s}$, which is equivalent with that $I_{d}:A_{u}^{s}\to L_{\mu}^{\frac{s(p+q^{\prime})}{pq^{\prime}}}$ is compact. Take $s=1$ then which implies that both $I_{d1}:A_{u}^{p}\to L_{\mu}^{\frac{s(p+q^{\prime})}{q^{\prime}}}$ and $I_{d2}:A_{u}^{q^{\prime}}\to L_{\mu}^{\frac{s(p+q^{\prime})}{p}}$ are compact.
Let $\{f_{n}\}$ be a bounded sequence in $A^{p}_{u}$. Since $I_{d1}:A_{u}^{p}\to L_{\mu}^{\frac{s(p+q^{\prime})}{q^{\prime}}}$ is compact, there exists a subsequence $\{f_{n_{k}}\}_{k}\subset\{f_{n}\}$ and for $f\in A_{u}^{p}$ there holds
$\lim_{k\to\infty}\|f_{n_{k}}-f\|_{L_{\mu}^{\frac{p+q^{\prime}}{q^{\prime}}}}=0.$
First, we consider the case $q>1$, let $g\in A_{u}^{q^{\prime}}$, by Hölder’s inequality we have
$|\langle T_{\mu}(f_{n_{k}}-f),g\rangle_{A_{u}^{2}}|$ $\leq$ $\int_{\mathbb{D}}|(f_{n_{k}}-f)(z)g(z)|d_{\mu(z)}$
$\leq$ $\|f_{n_{k}}-f\|_{L_{\mu}^{\frac{p+q^{\prime}}{q^{\prime}}}}\|g\|_{L_{\mu}^{\frac{p+q^{\prime}}{p}}}$
$\leq$ $\|f_{n_{k}}-f\|_{L_{\mu}^{\frac{p+q^{\prime}}{q^{\prime}}}}\|I_{d}\|_{L_{\mu}^{\frac{p+q^{\prime}}{p}}}\|g\|_{A_{u}^{q^{\prime}}}.$
Then, by duality from Lemma 2.12, we have
$\|T_{\mu}(f_{n_{k}}-f)\|_{A_{u}^{q}}$ $\simeq$ $\sup_{|g\|_{A_{u}^{q^{\prime}}}\leq 1}|\langle T_{\mu}(f_{n_{k}}-f),g\rangle_{A_{u}^{2}}|$
$\leq$ $\|f_{n_{k}}-f\|_{L_{\mu}^{\frac{p+q^{\prime}}{q^{\prime}}}}\|I_{d}\|_{L_{\mu}^{\frac{p+q^{\prime}}{p}}},$
which implies that $\|T_{\mu}(f_{n_{k}}-f)\|_{A_{u}^{q}}\to 0$ as $|z|\to 1^{-}$, then $T_{\mu}$ is compact.
[…]
$\|T_{\mu}(f_{n_{k}}-f)\|_{A_{u}^{q}}\lesssim\|f_{n_{k}}-f\|_{A_{u}^{q}},$
which implies that $\|T_{\mu}(f_{n_{k}}-f)\|_{A_{u}^{q}}\to 0$ as $|z|\to 1^{-}$, then $T_{\mu}$ is compact. ∎
5. Proof of Theorem 1.3
[…]
Suppose that Toeplitz operator $T_{\mu}:A^{p}_{u}\to A^{q}_{u}$ is bounded. For any arbitrary sequence $\lambda=\{\lambda_{k}\}\in\ell^{p}$, consider the function
$G_{t}(z)=\sum_{k=0}^{\infty}\lambda_{k}r_{k}(t)u(\Delta(z_{k},r))^{\frac{-1}{p}}K_{z_{k}}(z),\quad 0<t<1,$
where $r_{k}(t)$ is a sequence of Rademacher functions on $[0,1]$ and $\{z_{k}\}$ is a lattice constructed as Lemma 2.3. By estimate $\|K_{w}\|_{A_{u}^{p}}\simeq u(\Delta(w,r))^{\frac{1}{p}}/(1-|w|)^{2}$ in Lemma 2.6, we obtain
$\|G_{t}\|_{A^{p}_{u}}\lesssim\left(\sum_{k=0}^{\infty}|\lambda_{k}|^{p}\right)^{\frac{1}{p}}.$
Since $T_{\mu}:A^{p}_{u}\to A^{q}_{u}$ is bounded, gives
$\|T_{\mu}G_{t}\|_{A^{q}_{u}}^{q}\lesssim\|T_{\mu}\|^{q}\|G_{t}\|_{A_{u}^{p}}\lesssim\|T_{\mu}\|^{q}\|\lambda\|_{\ell^{p}}^{q},$
which implies that
$\int_{\mathbb{D}}\left|\sum_{k=0}^{\infty}\lambda_{k}r_{k}(t)u(\Delta(z_{k},r))^{\frac{-1}{p}}T_{\mu}K_{z_{k}}(z)\right|^{q}u(z)dA(z)\lesssim\|T_{\mu}\|^{q}\|\lambda\|_{\ell^{p}}^{q}.$
Integrating with respect to $t$ from 0 to 1, and by Fubini’s theorem and Khinchine’s inequality, we obtain
$B:=\int_{\mathbb{D}}\left(\sum_{k=0}^{\infty}|\lambda_{k}|^{2}u(\Delta(z_{k},r))^{\frac{-1}{p}}|T_{\mu}K_{z_{k}}(z)|^{2}\right)^{\frac{q}{2}}u(z)dA(z)\lesssim\|T_{\mu}\|^{q}\|\lambda\|_{\ell^{p}}^{q},$
let $\chi_{k}$ denote the characteristic function of $\Delta(z_{k},2r)$. By Lemma 2.3, the covering $\{\Delta(z_{k},2r)\}$ of $\mathbb{D}$ has finite multiply $N$,
$\sum_{k=0}^{\infty}|\lambda_{k}|^{q}u(\Delta(z_{k},r))^{\frac{-q}{2p}}\int_{\Delta(z_{k},2r)}|T_{\mu}K_{z_{k}}(z_{k})|^{q}u(z)dA(z)$
$=$ $\int_{\mathbb{D}}|\lambda_{k}|^{q}u(\Delta(z_{k},r))^{\frac{-q}{2p}}|T_{\mu}K_{z_{k}}|^{q}\chi_{k}(z)u(z)dA(z)$
$\leq$ $\max\{1,N^{1-\frac{q}{2}}\}B.$
Then Lemma 2.1 yields
(5.1) $\sum_{k=0}^{\infty}|\lambda_{k}|^{q}u(\Delta(z_{k},r))^{\frac{-q}{2p}+1}|T_{\mu}K_{z_{k}}(z_{k})|^{q}\lesssim\|T_{\mu}\|^{q}\|\lambda\|_{\ell^{p}}^{q}.$
On the other side, by kernel estimates obtained in Lemma 2.8, we have $|K_{z_{k}}(z)|^{2}\simeq\|K_{z_{k}}\|_{A_{u}^{2}}^{2}K(z,z)$, applying Lemma 2.6 and Lemma 2.8, we then obtain
$|T_{\mu}K_{z_{k}}(z_{k})|$ $\geq$ $\int_{\Delta(z_{k},r)}|K_{z_{k}}(z)|^{2}d\mu(z)$
$\simeq$ $\|K_{z_{k}}\|_{A_{u}^{2}}^{2}\int_{\Delta(z_{k},r)}K(z,z)d\mu(z)$
$\simeq$ $\frac{\widehat{\mu_{r}}(z_{k})u(\Delta(z_{k},r))}{(1-|z_{k}|)^{4}}$
$\geq$ $\widehat{\mu_{r}}(z_{k})u(\Delta(z_{k},r)).$
Therefore
$|T_{\mu}K_{z_{k}}(z_{k})|^{q}\gtrsim\widehat{\mu_{r}}(z_{k})^{q}u(\Delta(z_{k},r))^{q}.$
Then, (5.1) yields that
$\sum_{k=0}^{\infty}|\lambda_{k}|^{q}u(\Delta(z_{k},r))^{\frac{-q}{2p}+1+q}\widehat{\mu_{r}}(z_{k})^{q}\lesssim\|T_{\mu}\|^{q}\|\lambda\|_{\ell^{p}}^{q}.$
When $q<p$, we have $0<\frac{p-q}{p}<1+q(1-\frac{1}{2p})$,
$\sum_{k=0}^{\infty}|\lambda_{k}|^{q}u(\Delta(z_{k},r))^{\frac{p-q}{p}}\widehat{\mu_{r}}(z_{k})^{q}\lesssim\|T_{\mu}\|^{q}\|\lambda\|_{\ell^{p}}^{q}.$
By the duality between $\ell^{\frac{p}{q}}$ and $\ell^{\frac{p}{p-q}}$, we conclude that
$\sum_{k=0}^{\infty}\left(u(\Delta(z_{k},r))^{\frac{p-q}{p}}\widehat{\mu_{r}}(z_{k})^{q}\right)^{\frac{p}{p-q}}\lesssim\|T_{\mu}\|^{\frac{pq}{p-q}}.$
Thus
$\|\widehat{\mu_{r}}\|_{L^{\frac{pq}{p-q}}}^{\frac{pq}{p-q}}$ $=$ $\int_{\mathbb{D}}\widehat{\mu_{r}}(z)^{\frac{pq}{p-q}}dA(z)$
$\lesssim$ $\sum_{k=0}^{\infty}\int_{u(\Delta(z_{k},r))}\widehat{\mu_{r}}(z)^{\frac{pq}{p-q}}dA(z)$
$\lesssim$ $\sum_{k=0}^{\infty}u(\Delta(z_{k},r))\widehat{\mu_{r}}(z_{k})^{\frac{pq}{p-q}}$
$\lesssim$ $\|T_{\mu}\|^{\frac{pq}{p-q}}.$
which implies (iv).
[…]
Note that $p+1-\frac{p}{q}=\frac{p(p+q^{\prime})}{pq^{\prime}}$. Suppose that $\mu$ is a $\frac{p(p+q^{\prime})}{pq^{\prime}}$-vanishing Carleson measure for $A_{u}^{p}$, which is equivalent with that $I_{d}:A_{u}^{p}\to L_{\mu}^{\frac{p(p+q^{\prime})}{pq^{\prime}}}$ is compact. which implies that both $I_{d1}:A_{u}^{p}\to L_{\mu}^{\frac{(p+q^{\prime})}{q^{\prime}}}$ and $I_{d2}:A_{u}^{q^{\prime}}\to L_{\mu}^{\frac{(p+q^{\prime})}{p}}$ are compact.
Let $\{f_{n}\}$ be a bounded sequence in $A^{p}_{u}$. Since $I_{d1}:A_{u}^{p}\to L_{\mu}^{\frac{(p+q^{\prime})}{q^{\prime}}}$ is compact, there exists a subsequence $\{f_{n_{k}}\}_{k}\subset\{f_{n}\}$ and for $f\in A_{u}^{p}$ there holds,
$\lim_{k\to\infty}\|f_{n_{k}}-f\|_{L_{\mu}^{\frac{p+q^{\prime}}{q^{\prime}}}}=0.$
For $g\in A_{u}^{q^{\prime}}$, by Hölder’s inequality, we obtain,
$|\langle T_{\mu}(f_{n_{k}}-f),g\rangle_{A_{u}^{2}}|$ $\leq$ $\int_{\mathbb{D}}|(f_{n_{k}}-f)(z)g(z)|d{\mu(z)}$
$\leq$ $\|f_{n_{k}}-f\|_{L_{\mu}^{\frac{p+q^{\prime}}{q^{\prime}}}}\|g\|_{L_{\mu}^{\frac{p+q^{\prime}}{p}}}$
$\leq$ $\|f_{n_{k}}-f\|_{L_{\mu}^{\frac{p+q^{\prime}}{q^{\prime}}}}\|I_{d}\|_{L_{\mu}^{\frac{p+q^{\prime}}{p}}}\|g\|_{A_{u}^{q^{\prime}}}.$
Applying duality from Lemma 2.12, we obtain,
$\|T_{\mu}(f_{n_{k}}-f)\|_{A_{u}^{q}}$ $\simeq$ $\sup_{|g\|_{A_{u}^{q^{\prime}}}\leq 1}|\langle T_{\mu}(f_{n_{k}}-f),g\rangle_{A_{u}^{2}}|$
$\leq$ $\|f_{n_{k}}-f\|_{L_{\mu}^{\frac{p+q^{\prime}}{q^{\prime}}}}\|I_{d}\|_{L_{\mu}^{\frac{p+q^{\prime}}{p}}},$
which implies that $\|T_{\mu}(f_{n_{k}}-f)\|_{A_{u}^{q}}\to 0$ as $|z|\to 1^{-}$, and $T_{\mu}$ is compact.
[…]
For convenience, let $s=s(p,q)=p+1-\frac{p}{q}$, note that $\frac{pq}{pq+q-p}>1$ when $s>1$. By the subharmonicity of $|f|^{s}$, Lemma 2.1, Fubini’s theorem and Hölder’s inequality, we obtain,
$\int_{\mathbb{D}}|f(z)|^{s}d\mu(z)$ $\lesssim$ $\int_{\mathbb{D}}\frac{\int_{\Delta(z,r)}|f(\xi)|^{s}u(\xi)dA(\xi)}{u(\Delta(z,r))}d\mu(z)$
$\simeq$ $\int_{\mathbb{D}}\widehat{\mu_{r}}(\xi)|f(\xi)|^{s}u(\xi)dA(\xi)\leq\|f\|^{s}_{A_{u}^{p}}\|\widehat{\mu_{r}}\|_{L_{\mu}^{\frac{pq}{p-q}}}.$
6. Proof of Theorem 1.4 and Theorem 1.5
First, we recall some definitions about essential norm. Let $\mathcal{K}$ be the set of all compact operators on a Banach space $\mathcal{B}$. For any bounded linear operator $T:\mathcal{B}\to\mathcal{B}$, the essential norm of $T$ is defined by
$\|T\|_{e}=\inf\{\|T-K\|,K\in\mathcal{K}\}.$
And $\|T\|_{e}=0$ if and only if $T\in\mathcal{K}$.
Now we give the proof of Theorem 1.4.
Proof.
When $0<p\leq q<\infty$, by Lemma 2.10, it suffices to prove that,
$\limsup_{|z|\to 1}\widetilde{\mu_{t}}(z)u(\Delta(z,r))^{\frac{q-p}{pq}}\lesssim\|T_{\mu}\|_{e}\lesssim\limsup_{|z|\to 1}\widetilde{\mu_{t}}(z)u(\Delta(z,r))^{\frac{q-p}{pq}}.$
[…]
$\frac{|\widetilde{\mu_{t}}(z)|}{u(\Delta(z,r))^{\frac{1}{p}-\frac{1}{q}}}\lesssim\|T_{\mu}k_{p,z}\|_{A_{u}^{q}}.$
For any compact operator $K:{A^{p}_{u}\to A^{q}_{u}}$, using proposition 2.15, we obtain,
$\|Kk_{p,z}\|_{A_{u}^{q}}\to 0\text{ as }|z|\to 1^{-}$
Then
$\|T_{\mu}-K\|$ $\geq$ $\limsup_{|z|\to 1^{-}}\|(T_{\mu}-K)k_{p,z}\|_{A^{q}_{u}}$
$\geq$ $\limsup_{|z|\to 1^{-}}\|T_{\mu}k_{p,z}\|_{A^{q}_{u}}$
$\gtrsim$ $\limsup_{|z|\to 1^{-}}\widetilde{\mu_{t}}(z)u(\Delta(z,r))^{\frac{q-p}{pq}}$
Since $K$ is arbitrary, we have,
$\|T_{\mu}\|_{e}\gtrsim\limsup_{|z|\to 1^{-}}\widetilde{\mu_{t}}(z)u(\Delta(z,r))^{\frac{q-p}{pq}}.$
For the upper estimate, by Theorem 1.2, $T_{\mu_{0}}:A^{p}_{u}\to A^{q}_{u}$ is compact when $\mu_{0}$ is defined by assertions in Theorem 1.2. Using the norm equivalence from 1.1, we obtain,
$\|T_{\mu}-T_{\mu_{0}}\|\simeq\|\widehat{(\mu-\mu_{0})_{r}}u(\Delta(z,r))^{\frac{q-p}{pq}}\|_{L^{\infty}}=\sup_{z\in\mathbb{D}}\widehat{\mu_{r}}(z)u(\Delta(z,r))^{\frac{q-p}{pq}}$
Let $|z|\to 1^{-}$
$\|T_{\mu}\|\leq\|T_{\mu}-T_{\mu_{0}}\|\simeq\limsup_{|z|\to 1^{-}}\widehat{\mu_{r}}(z)u(\Delta(z,r))^{\frac{q-p}{pq}}.$
And we note that when $0<q<p<\infty$, by Theorem 1.3, $T_{\mu}$ is bounded if and only if $T_{\mu}$ is compact. Consequently, the essential norm of $T_{\mu}$ is 0.
Let $T$ be a compact operator on a Hilbert space $H$, and let $T^{*}$ be the Hilbert adjoint of $T$. By the spectral theorem, there exist orthonormal bases $\{\varphi_{k}\}$ and $\{\psi_{k}\}$ such that for $x\in H$,
$Tx=\sum_{k=1}^{\infty}\lambda_{k}\langle x,\varphi_{k}\rangle\psi_{k},$
where $\lambda_{k}$s are singular values of $T$. For $0<p<\infty$, the Schatten class $S_{p}=S_{p}(H)$ consists of those compact operators $T$ on $H$ whose singular value sequences $\{\lambda_{n}\}\in\ell^{p}$. For $P\geq 1$, the class $S_{p}$ forms a Banach space with the norm $\|T\|_{p}=\left(\sum_{n}|\lambda_{n}|^{p}\right)^{1/p}$.
Let $T$ be a compact operator and $h:\mathbb{R}^{+}\to\mathbb{R}^{+}$ be a continuous increasing function. We say that $T\in S_{h}(H)$ if there exists a positive constant $C>0$ such that
$\sum_{n=1}^{\infty}h(C\lambda_{n}(T))<\infty.$
Then, we give the proof of Theorem 1.5.
Proof.
Suppose that $T_{\mu}\in S_{h}(A_{u}^{2})$ . Let $\{e_{k}\}$ be an orthonormal set for $A^{2}_{u}$ and $T_{\mu}=\sum_{k=1}\lambda_{k}\langle\cdot,e_{k}\rangle e_{k}$ be the spectral decomposition of the positive operator $T_{\mu}$, where $\lambda_{k}$s are the eigenvalues of $T_{\mu}$. Since $\{e_{k}\}$ forms an orthonormal basis, we have $\sum_{k}|\langle k_{z},e_{k}\rangle_{A_{u}^{2}}|^{2}=1$ by Parseval Theorem. By Jensen’s formula, reproducing formula and the kernel estimate from Lemma 2.6, we obtain,
$\int_{\mathbb{C}^{n}}h(C\widetilde{\mu}_{2}(z))u(\Delta(z,r))^{-1}u(z)dA(z)$ $=$ $\int_{\mathbb{C}^{n}}h(C\langle T_{\mu}k_{z},k_{z}\rangle_{A_{u}^{2}})\frac{u(\Delta(z,r))}{(1-|z|)^{4}}u(z)dA(z)$
$=$ $\int_{\mathbb{C}^{n}}h\left(\sum_{k=1}^{\infty}C\lambda_{k}\left|\langle k_{z},e_{k}\rangle_{A_{u}^{2}}\right|^{2}\right)\frac{u(\Delta(z,r))}{(1-|z|)^{4}}u(z)dA(z)$
$\leq$ $\int_{\mathbb{C}^{n}}\sum_{k=1}^{\infty}h(C\lambda_{k})\left|\langle k_{z},e_{k}\rangle_{A_{u}^{2}}\right|^{2}\frac{u(\Delta(z,r))}{(1-|z|)^{4}}u(z)dA(z)$
$=$ $\int_{\mathbb{C}^{n}}\sum_{k=1}^{\infty}h(C\lambda_{k})\|K_{z}\|_{A^{2}_{u}}^{-2}\left|e_{k}(z)\right|^{2}\frac{u(\Delta(z,r))}{(1-|z|)^{4}}u(z)dA(z)$
$\simeq$ $\sum_{k=1}^{\infty}h(C\lambda_{k})\int_{\mathbb{C}^{n}}\left|e_{k}(z)\right|^{2}u(z)dA(z)$
$=$ $\sum_{k=1}^{\infty}h(C\lambda_{k})<\infty.$
Conversely, we assume that $\int_{\mathbb{C}^{n}}h(C\widetilde{\mu}_{2}(z))\frac{u(\Delta(z,r))}{(1-|z|)^{4}}u(z)dA(z)<\infty$ for some constant $C>0$. By applying Lemma 2.1, Fubini’s Theorem and Lemma 2.9, we obtain
$\langle T_{\mu}e_{k},e_{k}\rangle_{A_{u}^{2}}$ $=$ $\int_{\mathbb{C}^{n}}|e_{k}(z)|^{2}d\mu(z)$
$\lesssim$ $\int_{\mathbb{C}^{n}}\widehat{\mu}_{r}(z)|e_{k}(z)|^{2}u(z)dA(z)$
$\simeq$ $\int_{\mathbb{C}^{n}}\widetilde{\mu}_{2}(z)|e_{k}(z)|^{2}u(z)dA(z),$
Then, using Jensen’s formula again, we obtain,
$\sum_{k=1}^{\infty}h(C\langle T_{\mu}e_{k},e_{k}\rangle_{A_{u}^{2}})$ $\lesssim$ $\int_{\mathbb{C}^{n}}h(C\widetilde{\mu}_{2}(z))\left(\sum_{k\geq 1}|e_{k}(z)|^{2}\right)u(z)dA(z)$
$=$ $\int_{\mathbb{C}^{n}}h(C\widetilde{\mu}_{2}(z))\|K_{z}\|_{A_{u}^{2}}^{2}u(z)dA(z)$
$\simeq$ $\int_{\mathbb{C}^{n}}h(C\widetilde{\mu}_{2}(z))\frac{u(\Delta(z,r))}{(1-|z|)^{4}}u(z)dA(z)<\infty,$
which implies $T_{\mu}\in S_{h}(A_{u}^{2})$. ∎
Taking $h(t)=t^{p}$ in Theorem 1.5, we can get the following corollary which characterize Schatten class of Toeplitz operators on $A_{u}^{2}$.
Corollary 6.1.
Let $1\leq p<\infty$ and $\mu$ a positive Borel measure on $\mathbb{C}^{n}$ such that the Toeplitz operator $T_{\mu}:A_{u}^{2}\to A_{u}^{2}$ is compact. Then $T_{\mu}\in S_{p}(A_{u}^{2})$ if and only if
$\widetilde{\mu}_{2}(z)\left(\frac{u(\Delta(z,r))}{(1-|z|)^{4}}\right)^{1/p}\in A_{u}^{p}.$